# Higher equations of motion at level 2 in Liouville CFT

### v1

@@ 3. Boundary HEM
In this section, we prove the boundary version of the HEMs (Theorem 1.3). The introductory Section 3.1 gives some background on boundary LCFT. The remaining sections follow the structure of the proof of Theorem 1.1. All the technical difficulties are already present in the bulk case, and the differences are mostly computational. We will use most of the notations of Section 2, but all the notations from this section refer to boundary LCFT.
@@ 3.1.1. Free field modules
[…]
@@
[…]
@@
$\mathbb{E}[\mathbb{P}\varphi(z)P\varphi(w)]=\log\frac{1}{|1-z\bar{w}|}+\log\frac{1}{|1-zw|}=:G_{\partial}(z,w),\qquad\forall z,w\in\mathbb{D}\cap\mathbb{H}.$
@@
where $\mathrm{d}c$ is Lebesgue measure on $\mathbb{R}$. Samples of $\mathrm{d}c\otimes\mathbb{P}_{\mathbb{S}^{1}}$ are written $c+\varphi$, with $c$ being the zero mode of the field. We define a dense subspace $\mathcal{C}$ of $\mathcal{H}$ as the subspace of functionals $F\in\mathcal{H}$ such that there exists $N\in\mathbb{N}^{*}$ and $f\in\mathcal{C}^{\infty}(\mathbb{R}^{N+1})$, such that $F(c+\varphi)=f(c,\varphi_{1},...,\varphi_{N})$ holds $\mathrm{d}c\otimes\mathbb{P}_{\mathbb{S}^{1}}$-a.e., and $f$ and all its derivatives are compactly supported in $c$ and have at most exponential growth in the other variables. We refer to $\mathcal{C}$ as the space of test functions, and its continuous dual $\mathcal{C}^{\prime}$ as the space of tempered distributions.
Let $\alpha\in\mathbb{C}$. On $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$, we have a representation of the Heisenberg algebra $(\mathbf{A}_{n},\widetilde{\mathbf{A}}_{n})_{n\in\mathbb{Z}}$ given for $n>0$ by
@@
Here, $\partial_{n}=\partial_{\varphi_{n}}$ means (real) derivative in direction $\varphi_{n}$. For $n\neq 0$, we have the hermiticity relations $(\mathbf{A}_{n}^{\alpha})^{*}=\mathbf{A}_{-n}^{\alpha}$ on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. Given a partition $\mathbf{k}\in\mathcal{T}$, we set $\mathbf{A}_{-\mathbf{k}}:=\prod_{n=1}^{\infty}(\mathbf{A}_{-n}^{\alpha})^{k_{n}}$, and for $\mathbf{k}\in\mathcal{T}$, we set
$\pi_{\mathbf{k}}:=\mathbf{A}_{-\mathbf{k}}\mathds{1}.$
We say that $\pi_{\mathbf{k},\widetilde{\mathbf{k}}}$ has level $|\mathbf{k}|$. The Heisenberg representation gives $\mathcal{F}$ a structure of highest-weight Heisenberg module, i.e.
$\mathcal{F}=\mathrm{span}\,\{\pi_{\mathbf{k}}|\,\mathbf{k}\in\mathcal{T}\}.$
The level gives a grading
$\mathcal{F}=\oplus_{N\in\mathbb{N}}\mathcal{F}_{N},$
and $\dim\mathcal{F}_{N}=p(N)$.
@@
This representation satisfies the hermiticity relations $(\mathbf{L}_{n}^{0,\alpha})^{*}=\mathbf{L}_{-n}^{0,2Q-\bar{\alpha}}$ on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. Given a partition $\nu=(\nu_{1},...,\nu_{\ell})$, we set $\mathbf{L}_{-\nu}^{0,\alpha}=\mathbf{L}_{-\nu_{\ell}}^{0,\alpha}...\mathbf{L}_{-\nu_{1}}^{0,\alpha}$. The descendant state
@@
$\mathcal{V}_{\alpha}^{0}:=\mathrm{span}\{\mathbf{L}_{-\nu}^{0,\alpha}\mathds{1}|,\nu,\in\mathcal{T}\}\subset\mathcal{F}.$
We also define $\mathcal{V}_{\alpha}^{0,N}:=\mathcal{V}_{\alpha}^{0}\cap\mathcal{F}_{N}$. The module $\mathcal{V}_{\alpha}^{0}$ has central charge $c_{\mathrm{L}}=1+6Q^{2}$ and highest-weight $\Delta_{\alpha}=\frac{\alpha}{2}(Q-\frac{\alpha}{2})$. If $\alpha\not\in kac$, it is known that $\mathcal{V}_{\alpha}^{0}$ is irreducible and Verma (hence $\mathcal{V}_{\alpha}^{0}\simeq\mathcal{F}$) when $\alpha\not\in kac$. On the other hand, if $\alpha\in kac^{-}$, $\mathcal{V}_{2Q-\alpha}^{0}$ is Verma (hence $\mathcal{V}_{2Q-\alpha}^{0}\simeq\mathcal{F}$), and $\mathcal{V}_{\alpha}^{0}$ is the irreducible quotient of the Verma by the maximal proper submodule. The linear map
@@ 3.1.2. Semigroups and Poisson operator
Let $X_{\mathbb{D}}$ be a Dirichlet free field in $\mathbb{D}\cap\mathbb{H}$, i.e. $X_{\mathbb{D}}$ is Gaussian with covariance
@@
$\mathrm{d}L_{\gamma}(x):=\underset{\epsilon\to 0}{\lim}\,\epsilon^{\frac{\gamma^{2}}{4}}e^{\frac{\gamma}{2}X_{\epsilon}(x)}\mathrm{d}x.$
We will sometimes abuse notations by writing $\mathrm{d}M_{\gamma}(z)=e^{\gamma X(z)-\frac{\gamma^{2}}{2}\mathbb{E}[X^{2}(z)]}\frac{|\mathrm{d}z|^{2}}{(2\mathrm{Im}(z))^{\frac{\gamma^{2}}{2}}}$.
As in the bulk case, one can define two one-parameter semigroups of operators on $\mathcal{H}$. The free field semigroup is the semigroup generated by $-\mathbf{L}_{0}^{0}$. Currently, the generator of the Liouville semigroup is well-understood only if $\gamma<\sqrt{2}$ or $\mu=0$ [GRVW], but the semigroup itself is well-defined for all values of the parameters. Explictly, these two semigroups have the following expression:
@@
$S_{t}F(c+\varphi)=e^{-\frac{Q^{2}}{4}t}\mathbb{E}_{\varphi}\left[F(c+X(e^{-t}\cdot\,)e^{-\mu e^{\gamma c}\int_{\mathbb{A}_{t}\cap\mathbb{H}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma Q}}}e^{-\mu_{\mathrm{L}}e^{\frac{\gamma}{2}c}\int_{\mathbb{I}_{t}^{-}}\frac{\mathrm{d}L_{\gamma}(x)}{|x|^{\frac{\gamma Q}{2}}}}e^{-\mu_{\mathrm{R}}e^{\frac{\gamma}{2}c}\int_{\mathbb{I}_{t}^{+}}\frac{\mathrm{d}L_{\gamma}(x)}{|x|^{\frac{\gamma Q}{2}}}}\right],$
@@
We wish to construct a Poisson operator in a similar way as in the bulk case. For $\mathrm{Re}(\alpha)<Q$ and any $\chi\in\mathcal{F}$, we define (if the limit exists in $e^{-\beta c}\mathcal{H}$)
$\mathcal{P}_{\alpha}(\chi):=\underset{t\to\infty}{\lim}\,e^{t(\Delta_{\alpha}+|\nu|)}S_{t}(e^{\frac{1}{2}(\alpha-Q)c}\chi)$ (3.2)
$\Phi_{\alpha}(\chi):=\mathcal{P}_{\alpha}(\Phi_{\alpha}^{0}(\chi)).$
It is easy to see that the limit exists for $\alpha$ in a complex neighbourhood of $\infty$- ∞, but it is not known in full generality how the analytic extension behaves. This is actually the point of the HEMs to study the possible poles on the Kac table. Similar to the bulk case, it is expected that $\Psi_{\alpha,\nu}^{\partial}=\mathbf{L}_{-\nu}\Psi_{\alpha}^{\partial}$, for some Virasoro representation $(\mathbf{L}_{n})_{n\in\mathbb{Z}}$ defined in a similar fashion (and with similar properties) as the bulk representation [${}^{+}$23]. For the empty partition $\nu=\eset$, one has the expression (valid in a complex neighbourhood of $(-\infty,Q)$):
$\Psi_{\alpha}^{\partial}=e^{\frac{1}{2}(\alpha-Q)c}\mathbb{E}_{\varphi}\left[e^{-\mu\int_{\mathbb{D}\cap\mathbb{H}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}}e^{-\int_{\mathbb{I}}\mu_{\partial}(x)\frac{\mathrm{d}L_{\gamma}(x)}{|x|^{\frac{\gamma\alpha}{2}}}}\right],$
@@
Finally, we make the following definitions, with notations similar to the bulk case. For $\boldsymbol{w}\in(\mathbb{D}\cap\mathbb{H})^{r}$, $\boldsymbol{x}\in\mathbb{I}^{r_{\partial}}$, we introduce the following element of $e^{-\beta c}\mathcal{H}$:
@@
×$\qquad\times\left.\exp\left(-\mu e^{\gamma c}\int_{\mathbb{D}\cap\mathbb{H}}\epsilon^{\frac{\gamma^{2}}{2}}e^{\gamma X_{\epsilon}(z)}|\mathrm{d}z|^{2}-e^{\frac{\gamma}{2}c}\int_{\mathbb{I}}\epsilon^{\frac{\gamma^{2}}{4}}e^{\frac{\gamma}{2}X_{\epsilon}(x)}\mu_{\partial}(x)\mathrm{d}x\right)\right],$
[…]
$\mathcal{I}^{\partial}_{\boldsymbol{s},\widetilde{\boldsymbol{s}}}(\alpha):=\int_{\mathbb{I}_{-}^{r}}\int_{\mathbb{I}_{+}^{r^{\prime}}}\Psi_{\alpha}^{\partial}(x_{1},...,x_{r},\widetilde{x}_{1},...,\widetilde{x}_{\widetilde{r}})\prod_{j=1}^{r}\frac{\mathrm{d}x_{j}}{|x_{j}|^{s_{j}}}\prod_{\widetilde{j}=1}^{\widetilde{r}}\frac{\mathrm{d}\widetilde{x}_{\widetilde{j}}}{|\widetilde{x}_{\widetilde{j}}|^{\widetilde{s}_{j}}}.$ (3.3)
[…]
@@
In this work, we only focus on the singular vector at level 2, whose free field expression is given in Lemma 3.1. Namely, we are interested in the analytic continuation of $\Phi_{\alpha}(\mathbf{S}^{0}_{2Q-\alpha}\mathds{1})$, particularly its value at $\alpha_{1,2},\alpha_{2,1}$. We keep the notation $\Phi_{\alpha}(\mathbf{S}_{2Q-\alpha}^{0}\mathds{1})$ for this analytic continuation, wherever it is defined. The remainder of this section is devoted to the proof of Theorem 1.3, following the strategy of Section 2. First, we find an probabilistic expression for the singular state, valid up to $\alpha_{1,2}$. This leads to the $(1,2)$-HEM without too much effort. Then, we find an expression for the meromorphic continuation of the singular state, valid up to $\alpha_{2,1}$. The $(2,1)$-HEM is obtained by evaluating the residue at $\alpha_{2,1}$.

### v2

@@ 3. Boundary HEM
In this section, we prove the boundary version of the HEMs (Theorem 1.2). This section and the notations therein are independent from Section 2; for instance, the space $\mathcal{F}$ and measure $\mathbb{P}_{\mathbb{S}^{1}}$ introduced below are different from (although very similar to) those in Section 2. The introductory Section 3.1 gives some background on boundary LCFT. The remaining sections follow the structure of the proof of Theorem 1.1. All the technical difficulties are already present in the bulk case, and the differences are mostly computational. We will use most of the notations of Section 2, but all the notations from this section refer to boundary LCFT.
@@ 3.1.1. Free field modules
Let $\mathcal{F}:=\mathbb{R}[(\varphi_{n})_{n\;\geqslant\;1}]$ be the space of polynomials in countably many real variables $\varphi_{n}$. The constant function $\mathds{1}$ is the vacuum vector. Let $\mathbb{P}_{\mathbb{S}^{1}}$ be the law of a $\log$-correlated Gaussian field $\varphi$ on $\mathbb{S}^{1}\cap\mathbb{H}$:
@@
where we recall that $e_{n}(e^{i\theta})=e^{ni\theta}$ is the standard basis. Under $\mathbb{P}_{\mathbb{S}^{1}}$, $(\varphi_{n})_{n\;\geqslant\;1}$ is a sequence of independent real Gaussians $\mathcal{N}(0,\frac{1}{2n})$. The harmonic extension of $\varphi$ to $\mathbb{D}\cap\mathbb{H}$ is
@@
$\mathbb{E}[P\varphi(z)P\varphi(w)]=\log\frac{1}{|1-z\bar{w}|}+\log\frac{1}{|1-zw|}=:G_{\partial}(z,w),\qquad\forall z,w\in\mathbb{D}\cap\mathbb{H}.$
@@
where $\mathrm{d}c$ is Lebesgue measure on $\mathbb{R}$. Samples of $\mathrm{d}c\otimes\mathbb{P}_{\mathbb{S}^{1}}$ are written $c+\varphi$, with $c$ being the zero mode of the field. We define a dense subspace $\mathcal{C}$ of $\mathcal{H}$ as the subspace of functionals $F\in\mathcal{H}$ such that there exists $N\in\mathbb{N}^{*}$ and $f\in\mathcal{C}^{\infty}(\mathbb{R}^{N+1})$, such that $F(c+\varphi)=f(c,\varphi_{1},...,\varphi_{N})$ holds $\mathrm{d}c\otimes\mathbb{P}_{\mathbb{S}^{1}}$-a.e., and $f$ and all its derivatives are compactly supported in $c$ and have at most exponential growth in the other variables. The space $\mathcal{C}$ comes with a natural Fréchet topology, and we denote by $\mathcal{C}^{\prime}$ the space of continuous linear forms on $\mathcal{C}$.
Let $\alpha\in\mathbb{C}$. We have a representation of the Heisenberg algebra $(\mathbf{A}_{n}^{\alpha})_{n\in\mathbb{Z}}$ acting as densely defined operators on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$, given for $n>0$ by
@@
Here, $\partial_{n}=\partial_{\varphi_{n}}$ is the real derivative in the direction $\varphi_{n}$. For $n\neq 0$, we have the hermiticity relations $(\mathbf{A}_{n}^{\alpha})^{*}=\mathbf{A}_{-n}^{\alpha}$ on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$.
@@
This representation satisfies the hermiticity relations $(\mathbf{L}_{n}^{0,\alpha})^{*}=\mathbf{L}_{-n}^{0,2Q-\bar{\alpha}}$ on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. Given a Young diagram $\nu=(\nu_{1},...,\nu_{\ell})$, we set $\mathbf{L}_{-\nu}^{0,\alpha}=\mathbf{L}_{-\nu_{\ell}}^{0,\alpha}...\mathbf{L}_{-\nu_{1}}^{0,\alpha}$. The descendant state
@@
$\mathcal{V}_{\alpha}^{0}:=\mathrm{span}\{\mathbf{L}_{-\nu}^{0,\alpha}\mathds{1}|,\nu\in\mathcal{T}\}\subset\mathcal{F}.$
We also define $\mathcal{V}_{\alpha}^{0,N}:=\mathcal{V}_{\alpha}^{0}\cap\mathcal{F}_{N}$. The module $\mathcal{V}_{\alpha}^{0}$ has central charge $c_{\mathrm{L}}=1+6Q^{2}$ and highest-weight $\Delta_{\alpha}=\frac{\alpha}{2}(Q-\frac{\alpha}{2})$. If $\alpha\not\in kac$, it is known that $\mathcal{V}_{\alpha}^{0}$ is irreducible and Verma [KRR13, Lecture 8] (see also [BW23, Appendix B]), hence $\mathcal{V}_{\alpha}^{0}\simeq\mathcal{F}$). On the other hand, if $\alpha\in kac^{-}$, $\mathcal{V}_{2Q-\alpha}^{0}$ is Verma (hence $\mathcal{V}_{2Q-\alpha}^{0}\simeq\mathcal{F}$), and $\mathcal{V}_{\alpha}^{0}$ is the irreducible quotient of the Verma by the maximal proper submodule. The linear map
@@ 3.1.2. Semigroups and Poisson operator
Let $X_{\mathbb{D}}$ be the GFF in $\mathbb{H}\cap\mathbb{D}$ with Dirichlet boundary conditions on $\mathbb{S}^{1}\cap\mathbb{H}$ and Neumann boudary conditions on (1,1)(-1,1). Explicitly, $X_{\mathbb{D}}$ is the Gaussian field with covariance:
@@
$\mathrm{d}L_{\gamma}(x):=\underset{\epsilon\to 0}{\lim}\,\epsilon^{\frac{\gamma^{2}}{4}}e^{\frac{\gamma}{2}X_{\epsilon}(x)}\mathrm{d}x,$
where we use the same regularisation as in [GRW24b, Section 2.9]: we use circle averages in the bulk, and semi-circle averages on (1,1)(-1,1). We will sometimes abuse notations by writing $\mathrm{d}M_{\gamma}(z)=e^{\gamma X(z)-\frac{\gamma^{2}}{2}\mathbb{E}[X^{2}(z)]}\frac{|\mathrm{d}z|^{2}}{(2\mathrm{Im}(z))^{\frac{\gamma^{2}}{2}}}$.
As in the bulk case, one can define two one-parameter semigroups of operators on $\mathcal{H}$. The free field semigroup is the semigroup generated by $\mathbf{L}_{0}^{0}$. Currently, the generator of the Liouville semigroup is well-understood only if $\gamma<\sqrt{2}$ or $\mu=0$ [GRW24b], but the semigroup itself is well-defined for all values of the parameters. Explictly, these two semigroups have the following expression [GRW24b, Section 7.2] (see in particular Proposition 7.3 therein):
@@
$S_{t}F(c+\varphi)=e^{-\frac{Q^{2}}{4}t}\mathbb{E}_{\varphi}\left[F(c+X(e^{-t}\cdot\,))e^{-\mu e^{\gamma c}\int_{\mathbb{A}_{t}\cap\mathbb{H}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma Q}}}e^{-\mu_{\mathrm{L}}e^{\frac{\gamma}{2}c}\int_{\mathbb{I}_{t}^{-}}\frac{\mathrm{d}L_{\gamma}(x)}{|x|^{\frac{\gamma Q}{2}}}}e^{-\mu_{\mathrm{R}}e^{\frac{\gamma}{2}c}\int_{\mathbb{I}_{t}^{+}}\frac{\mathrm{d}L_{\gamma}(x)}{|x|^{\frac{\gamma Q}{2}}}}\right],$
@@
We wish to construct a Poisson operator in a similar way as in the bulk case. For $\mathrm{Re}(\alpha)<Q$ and any $\nu\in\mathcal{T}$, we define (if the limit exists in the weighted space $e^{-\beta c}\mathcal{H}$ for some $\beta\in\mathbb{R}$)
$\mathcal{P}_{\alpha}(\mathcal{Q}_{\alpha,\nu})=\underset{t\to\infty}{\lim}\,e^{t(\Delta_{\alpha}+|\nu|)}S_{t}(e^{\frac{1}{2}(\alpha-Q)c}\mathcal{Q}_{\alpha,\nu})$ (3.2)
$\Phi_{\alpha}(\chi)=\mathcal{P}_{\alpha}(\Phi_{\alpha}^{0}(\mathcal{Q}_{\alpha,\nu})).$
Applied to the constant function $\chi=1$ and for $\alpha<Q$, the limiting function is seen to be given by
$\Psi_{\alpha}^{\partial}:=\Phi_{\alpha}(\mathds{1})=e^{\frac{1}{2}(\alpha-Q)c}\mathbb{E}_{\varphi}\left[e^{-\mu\int_{\mathbb{D}\cap\mathbb{H}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}}e^{-\int_{\mathbb{I}}\mu_{\partial}(x)\frac{\mathrm{d}L_{\gamma}(x)}{|x|^{\frac{\gamma\alpha}{2}}}}\right],$
@@
The goal of this section is to study the state $\Phi_{\alpha}(\mathbf{S}_{2Q-\alpha}^{0}\mathds{1})$. We will find a probabilistic expression for $\alpha<\alpha_{1,2}$, and study its meromorphic continuation. As in the previous section, since we only work at level 2 and for $\alpha$ bounded below, we can fix once and for all a large $\beta>0$ such that all the states consider in this section belong to the weighted space $e^{-\beta c}\mathcal{H}$.
Now, we make some definitions. For $\boldsymbol{w}\in(\mathbb{D}\cap\mathbb{H})^{r}$, $\boldsymbol{x}\in\mathbb{I}^{r_{\partial}}$, we introduce the following element of $e^{-\beta c}\mathcal{H}$:
@@
$\qquad\times\left.\exp\left(-\mu e^{\gamma c}\int_{\mathbb{D}\cap\mathbb{H}}\epsilon^{\frac{\gamma^{2}}{2}}e^{\gamma X_{\epsilon}(z)}|\mathrm{d}z|^{2}-e^{\frac{\gamma}{2}c}\int_{\mathbb{I}}\epsilon^{\frac{\gamma^{2}}{4}}e^{\frac{\gamma}{2}X_{\epsilon}(x)}\mu_{\partial}(x)\mathrm{d}x\right)\right],$
In practice, we will only be dealing with cases where $2r+r_{\partial}\;\leqslant\;3$. Finally, given Young diagrams $\boldsymbol{s}=(s_{1},...,s_{r})\in\mathbb{N}^{r}$ and $\widetilde{\boldsymbol{s}}=(\widetilde{s}_{1},...,\widetilde{s}_{\widetilde{r}})\in\mathbb{N}^{\widetilde{r}}$ we introduce the integrals (provided they are well-defined in $e^{-\beta c}\mathcal{H}$)
$\mathcal{I}^{\partial}_{\boldsymbol{s},\widetilde{\boldsymbol{s}}}(\alpha):=\int_{\mathbb{I}_{-}^{r}}\int_{\mathbb{I}_{+}^{\widetilde{r}}}\Psi_{\alpha}^{\partial}(x_{1},...,x_{r},\widetilde{x}_{1},...,\widetilde{x}_{\widetilde{r}})\prod_{j=1}^{r}\frac{\mathrm{d}x_{j}}{|x_{j}|^{s_{j}}}\prod_{\widetilde{j}=1}^{\widetilde{r}}\frac{\mathrm{d}\widetilde{x}_{\widetilde{j}}}{|\widetilde{x}_{\widetilde{j}}|^{\widetilde{s}_{j}}}.$ (3.3)
In practice, we will only deal with cases where $|\boldsymbol{s}|+|\widetilde{\boldsymbol{s}}|\;\leqslant\;2$. In particular, the total number of insertions is bounded by 2. We also define
@@
Since this does not seem to have appeared in the literature yet, we include the following lemma on the analytic continuation of $\Psi_{\alpha}^{\partial}(\boldsymbol{w})$, which is a straightforward adaptation of [KRV20, Theorem 6.1].
Lemma 3.2.
In the setting above, for all $\beta>Q-\alpha-r\gamma$ and $\epsilon>0$, the map $\alpha\mapsto\Psi_{\alpha}^{\partial}(\boldsymbol{w})\in e^{-\beta c}\mathcal{H}$ has an analytic continuation in a complex neighbourhood of $(-\infty,Q-\epsilon)$.
Proof.
For notational simplicity and since the question is about the local behaviour of GMC around 0, we only treat the case with no $\gamma$-insertions at the other marked points (i.e. $r=0$). The strategy applies verbatim to the case $r\;\geqslant\;1$.
Let $\alpha\in(-\infty,Q-\epsilon)$ and define the a.s. increasing function of $t\;\geqslant\;0$:
$Z_{t}^{\alpha}:=\mu e^{\gamma c}\int_{\mathbb{A}_{t}\cap\mathbb{H}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}+e^{\frac{\gamma}{2}c}\int_{\mathbb{I}\setminus e^{-t}\mathbb{I}}\mu_{\partial}(x)\frac{\mathrm{d}L_{\gamma}(x)}{|x|^{\frac{\gamma\alpha}{2}}}.$
Under the free field semigroup (3.1), the zero mode evolves as an independent Brownian motion at speed 2: $c_{t}=c+B_{2t}$. (Compared to the bulk scenario, the speed 2 comes from the fact that the Neumann Green function in $\mathbb{H}$ is $\log\frac{1}{|z-w|}+\log\frac{1}{|z-\bar{w}|}$, which behaves as twice the $\log$ when $w\to 0$.) By the Cameron–Martin formula, reweighting by $e^{\frac{\alpha}{2}B_{2t}-\frac{\alpha^{2}}{4}t}$ amounts to shifting the field by $\alpha\max(\log|\frac{1}{z}|,t)$, and we therefore get $\Psi_{\alpha}^{\partial}=e^{\frac{1}{2}(\alpha-Q)c}\lim_{t\to\infty}\mathbb{E}_{\varphi}[e^{-Z_{t}^{\alpha}}]$ (which is finite for $\alpha<Q$ [GKR+24a]).
Let $F\in\mathcal{C}$, and for all $t>0$, set $\alpha\mapsto f_{t}(\alpha):=\mathbb{E}[e^{\frac{\alpha}{2}B_{2t}-\frac{\alpha^{2}}{4}t}e^{-Z_{t}^{0}}F]\in e^{-\beta c}\mathcal{H}$. The function $f_{t}$ is analytic in the whole $\alpha$-plane for each $t>0$, hence it suffices to show that $f_{t}$ converges uniformly on compacts of some neighbourhood of $(-\infty,Q-\epsilon)$, as $t\to\infty$. Arguing as in the proof of Proposition 2.6, it is sufficient to consider the case $F=1$. Following the proof of [KRV20, Theorem 6.1], and using the elementary inequality $1-e^{-x}\;\leqslant\;x^{p}$ for $x>0$ and $p\in(0,1)$, we get:
$|f_{t}(\alpha)-f_{t-1}(\alpha)|\;\leqslant\;e^{\mathrm{Im}(\alpha)^{2}\frac{t}{4}}\mathbb{E}[1-e^{-(Z_{t}^{\mathrm{Re}(\alpha)}-Z_{t-1}^{\mathrm{Re}(\alpha)})}]\;\leqslant\;e^{\mathrm{Im}(\alpha)^{2}\frac{t}{4}}\mathbb{E}[(Z_{t}^{\mathrm{Re}(\alpha)}-Z_{t-1}^{\mathrm{Re}(\alpha)})^{p}]=e^{(\frac{1}{4}\mathrm{Im}(\alpha)^{2}-\xi)t}\mathbb{E}[(Z_{1}^{\mathrm{Re}(\alpha)})^{p}].$
Here, $p>0$ is chosen small enough so that the last moment is finite [HRV18, (3.23)], and $\xi=\xi(p,\mathrm{Re}(\alpha))>0$ is a multifractal exponent of the type already encountered in the proof of Proposition 2.6. It is readily seen that $\xi$ is affine and non-increasing in the variable $\mathrm{Re}(\alpha)$, so that $f_{t}$ converges uniformly in the region $\{\alpha\in\mathbb{C}|\,\mathrm{Re}(\alpha)<Q-\epsilon\text{ and }|\mathrm{Im}(\alpha)|<2\sqrt{\xi(p,Q-\epsilon)}\}$. ∎